# Random Policy Enables In-Context Reinforcement Learning within Trust Horizons

### v2

@@ 1 Introduction
Pretrained foundation models (FMs) have demonstrated promising performance across a wide variety of domains in artificial intelligence including natural language processing (NLP) (Devlin, 2018; Radford, 2018; Radford et al., 2019; Brown, 2020), computer vision (CV) (Yuan et al., 2021; Sammani et al., 2022; Ma et al., 2023; Chen et al., 2024b), and sequential decision-making (Chen et al., 2021; Janner et al., 2021; Xu et al., 2022b; Yang et al., 2023; Light et al., 2024a; b). This success is attributed to FMs’ impressive capability of in-context learning (Dong et al., 2022; Li et al., 2023; Wei et al., 2023; Wies et al., 2024) which refers to the ability to infer and understand the new tasks provided with the context information (or prompt) and without model parameters updates. Recently, in-context reinforcement learning (ICRL) (Laskin et al., 2022; Grigsby et al., 2023; Lin et al., 2023; Sinii et al., 2023; Zisman et al., 2023; Lee et al., 2024; Lu et al., 2024; Wang et al., 2024; Dong et al., 2024) has emerged when FMs are pretrained on sequential decision-making problems. Whereas FMs use texts as the context/prompt in NLP, ICRL treats the state-action-reward tuples as the contextual information for decision-making.
@@
However, the current state-of-the-art (SOTA) ICRL algorithms impose strict requirements on the pretraining datasets. More specifically, Algorithm Distillation (AD) (Laskin et al., 2022) requires the context to contain the complete learning history (from the initial policy to the final-trained policy) of the source (or behavior) RL algorithm for all pretraining environments. In addition, AD requires environments to have short episodes, allowing the context to capture cross-episodic information. This enables AD to learn the improvement operator of the source RL algorithm. Conversely, Decision Pretrained Transformer (DPT) (Lee et al., 2024) partially relaxes the requirement on the context, permitting it to be gathered by random policies and without needing to adhere to the transition dynamics. Nevertheless, DPT necessitates the optimal policy to label an optimal action for any randomly sampled query state across all pretraining environments. To explore the feasibility of ICRL in the absence of optimal policies, Decision Importance Transformer (DIT) (Dong et al., 2024) proposes to leverage the observed state-action pairs in the context data as query states and corresponding action labels. Each state-action pair within the context is assigned a weight in the training process. This weight is proportional to the pseudo-return of the pair. Thus, DIT prioritizes the training on high-pseudo-return pairs. Despite not demanding optimal policies, DIT still requires a substantial context dataset to comprehensively cover all state-action pairs from the state and action spaces. Furthermore, DIT mandates that more than $30\%$ of the transition data in the context be well-trained, and the context should originate from a complete episode.
@@
Notably, acquiring either optimal policies or well-trained policies across a multitude of pretraining environments in real-world scenarios can be prohibitively expensive or even intractable. On the other hand, the transition data available in real-world problems like healthcare (Fatemi et al., 2022; Tang & Wiens, 2021)—collected as the context—may not originate from a complete episode. These stringent requirements on the pretraining dataset of the SOTA ICRL algorithms severely limit their practical applications to the real world, especially for those where the transition data exhibits high variance to train an effective policy, and a naive random policy such as the uniform policy becomes a reasonable choice at hand. Consequently, this paper centers on the ICRL that operates without the need for optimal (or any degree of well-trained) policies or episodic context, placing its emphasis on the scenarios under (e.g., uniform) random policies and random contexts only. More background knowledge can be found in the Related Work section (see Appendix A).
Figure 1: Schematic of the State-Action Distillation (SAD) approach. SAD first collects the context and query state using a random policy, then distills the action label by exploring the action space using the random policy and query state within a trust horizon. The collected contexts, query states, and action labels form the pretraining dataset, which can be employed to pretrain a foundation model for in-context reinforcement learning in a supervised pretraining mechanism.
@@ 2.1 Preliminaries
RL problems are generally formulated as Markov Decision Processes (MDPs) (Sutton, 2018). An MDP can be represented by a tuple $\tau=(S,A,R,P,\rho)$, where $S$ and $A$ denote finite state and action spaces, $R:S\times A\to\mathbb{R}$ denotes the reward function that evaluates the quality of the decision (action), $P:S\times A\times S\to[0,1]$ denotes the transition probability that describes the dynamics of the system, and $\rho:S\to[0,1]$ denotes the initial state distribution.
@@ 2.2 Supervised Pretraining Mechanism
In this subsection, we introduce the methodology behind ICRL–a supervised pretraining mechanism. Consider two distributions over environments ${\mathcal{T}}_{\text{train}}$ and ${\mathcal{T}}_{\text{test}}$ for pretraining and test (evaluation), respectively. Each environment, along with its corresponding MDP $\tau$, can be regarded as an instance drawn from the environment distributions, where each environment may exhibit distinct reward functions and transition dynamics. Given an environment $\tau$, a context/prompt ${\mathcal{C}}=\{s_{i},a_{i},r_{i},s_{i}^{\prime}\}_{i\in[n]}$ refers to a collection of interactions between the agent and the environment $\tau$, sampled from a pretraining context distribution $D_{\text{train}}(\cdot\mid\tau)$, i.e., ${\mathcal{C}}\sim D_{\text{train}}(\cdot\mid\tau)$. Notably, $D_{\text{train}}(\cdot\mid\tau)$ contains the contextual information regarding the environment $\tau$. We next consider a query state distribution $D_{q}^{\tau}$ and a label policy that maps the query state to the action label, i.e., $\pi_{l}:S\to\Delta(a_{l})$. Then, the joint distribution over the environment $\tau$, context ${\mathcal{C}}$, query state $s_{q}$, and action label $a_{l}$ is given by
@@
where $l(\cdot,\cdot)$ represents the loss function, for instance, negative log-likelihood (NLL) for discrete-action problems and mean square error (MSE) for continuous-action scenarios. We note that while the current SOTA ICRL algorithms (AD, DPT, and DIT) adhere to a common objective function equation 3, they differ significantly in constructing the context, query state, and action label.
@@
Instead of learning an optimal policy for a specific environment, AD proposes to learn a RL algorithm itself across a wide range of environments. This is, the FM in AD is pretrained to imitate the source (or behavior) algorithm over the pretraining environment distribution ${\mathcal{T}}_{\text{train}}$. In general, AD demands a well-trained source algorithm with its complete learning history (from the initial policy to the final-trained policy). Additionally, AD is restricted to the environments with short episode length by which the context must capture cross-episode information of the source algorithm. In terms of the objective function equation 3, AD takes the state at the time step $t$ as the query state, $a_{t}$ from the source algorithm as the action label, and the episodic history data $(s_{0},a_{0},r_{0},\cdots,s_{t-1},a_{t-1},r_{t-1})$ to be the context.
@@
Instead of being stringent on the pretraining context dataset and the environment itself, DPT handles the context and query state in a more general manner. Specifically, DPT considers an entirely random collection of transitions as the context, a random query state $s_{q}$ drawn from $D_{q}$, and an optimal action label corresponding to $s_{q}$. Despite less requirements on the context, DPT necessitates access to optimal policies for the optimal action labels in all pretraining environments, which may not be available in real-world applications.
@@
DIT proposes to learn ICRL without optimal action labels, following the same supervised pretraining mechanism as DPT. To that end, DIT leverages every possible state and action in the context as the query state and corresponding action label. It is important to point out that DIT requires a (partially) complete episode to form the context, enabling the computation of a pseudo-return for each state-action pair. By mapping the pseudo-return to a weight that reflects the quality of each state-action pair, DIT can pretrain the FM using the DPT structure, augmented by the weight assigned to each query state and action label. In other words, DIT prioritizes the training on high-pseudo-return pairs. Last but not least, DIT still mandates that more than $30\%$ of the context data comes from well-trained policies.
To summarize, it is worth highlighting that all these SOTA ICRL algorithms necessitate varying degrees of well-trained, or even optimal policies during the pretraining phase. However, obtaining optimal or well-trained policies for real-world applications is often prohibitive, as it demands extensive training across a vast number of real-world environments. This challenge becomes even more pronounced in the domains like healthcare (Fatemi et al., 2022), where the transition data exhibits high variance to train an effective policy, and the random policy becomes a reasonable choice at hand. Therefore, executing ICRL under (e.g., uniform) random policies and random contexts is crucial for enabling the practical application of ICRL in the real-world.
@@ 3 State-Action Distillation
In this section, we propose the State-Action Distillation (SAD), an approach for generating the pretraining dataset for ICRL under random policies and random contexts (see Figure 1). We summarize the implementation details of SAD in Algorithms 1-6.
As indicated in equation 2, the pretraining data consists of the context, the query state and the action label. We start by introducing the generation of the context under random policies in SAD (refer to Algorithm 1). It is important to highlight that the context is derived through interactions with the environment under any given random policy (e.g., uniform random policy), and notably, the context not necessarily originates from a complete episode. These benefits make SAD potentially well-suited for ICRL’s real-world applications with random transition data only.
@@
We proceed by recalling that DIT prioritizes the training on high-pseudo-return pairs from the context data collected. To that end, DIT assigns a weight $w$ to the loss function during the pretraining phase that is proportional to the pseudo-return, i.e., $w(s_{t},a_{t})\propto\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}.$ Nonetheless, we acknowledge that DIT may not explore to train on good state-action pairs under the random policy for two reasons: ($i$) DIT solely considers to train on the state-action pairs that are observed in the context, which contains limited transition data and is derived by the random policy. ($ii$) Even for the state-action pairs in the collected context, the pseudo-return does not necessarily prioritize the optimal pair but rather promotes the pair with high immediate reward, as the discount factor applies starting from the current time step with a horizon of $(T-t+1)$ only, instead of $T$.
Under this circumstance, our SAD approach advocates for distilling the outstanding query states and action labels by searching across the entire state and action spaces under the random policy. Before proceeding, we first recall the definition of the optimal action in the problems of multi-armed bandit (MAB) and MDP. For any query state $s_{q}$, the optimal action in the action space $A$ for $s_{q}$ corresponds to the action that maximizes the optimal Q-function, i.e.,
$a_{\text{MAB}}^{*}(s_{q})\overset{\Delta}{=}\operatorname*{arg\,max}_{a\in A}\underbrace{{\mathbb{E}}\left[r(s_{q},a)\right]}_{Q_{\text{MAB}}^{*}(s_{q},a)},$ (4)
$a_{\text{MDP}}^{*}(s_{q})\overset{\Delta}{=}\operatorname*{arg\,max}_{a\in A}\underbrace{{\mathbb{E}}_{\pi^{*}}\!\!\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}=s_{q},a_{0}=a\right]}_{Q_{\text{MDP}}^{*}(s_{q},a)},$ (5)
where $s_{q}$ in equation 4 refers to the singleton state in the MAB problem, and $\pi^{*}$ in equation 5 denotes the optimal policy in the MDP.
However, both $a_{\text{MAB}}^{*}$ and $a_{\text{MDP}}^{*}$ are intractable to obtain in our problem of interest for two reasons: ($i$) computing the expectation in the Q-function demands to sample infinite episodes; ($ii$) one can only have access to the random policy, instead of $\pi^{*}$. Therefore, we instead consider ($i$) stochastic approximation that uses the average as the unbiased estimate of the expectation due to the law of large numbers; ($ii$) maximizing the return under the random policy.
@@ 3.1 Trustworthiness of the Random Policy
Subsequently, the crucial question arises: when can we trust the random policy? We claim: The random policy is probabilistically trustworthy for the MAB and MDP problems within a trust horizon. We formalize this claim for the MAB and MDP respectively in this subsection, which relies on the following assumption.
@@
Note that Assumption 1 is common in the literature (Azar et al., 2017; Wei et al., 2020; Zhang et al., 2021). In particular, in the case of finite state-action spaces, it is always possible to design the reward to avoid the possibility of being unbounded. With Assumption 1 established, we first formalize the trustworthiness of the random policy for the MAB problem in the following theorem.
@@
Let Assumption 1 hold. Denote by $s_{q}$ and $a^{*}$ the singleton state and optimal arm in the MAB problem. Consider a random policy $\pi$. Suppose that each arm has been selected $N$ times ($N\in{\mathbb{N}}_{+}$, trust horizon) under $\pi$. With probability at least $1-\delta$, it holds that
$\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq\max_{a\in A\setminus\{a^{*}\}}\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi),$ (6)
when the trust horizon $N$ satisfies
@@
Proof. See Appendix B.1.
$\square$
Theorem 1 implies that the trust horizon $N$ quantifies the trustworthiness of the decision making under the random policy $\pi$ for MAB problems. Indeed, a larger $N$ implies a higher probability (smaller $\delta$) that the average reward of the optimal arm under $\pi$ exceeds that of the next-best arm, therefore, making a more reliable decision. We substantiate this claim by empirical evidence (depicted in Figure 5).
In the practical implementation, we simply execute the MAB under the random policy $\pi$ until every action in the action space $A$ selected exactly $N$ times, discarding any actions that exceed the trust horizon $N$. Subsequently, we select the action with the maximal average reward as the action label for the singleton state. The detailed procedure for collecting such action labels in MAB is outlined in Algorithm 2.
@@
2:  Execute the MAB in $\tau$ under the random policy $\pi$ until every action in $A$ selected exactly $N$ times, discarding any actions that exceed the trust horizon $N$.
@@
While the trustworthiness of the random policy in the MAB problem has been established and discussed, the scenario of the MDP presents a distinct challenge. To proceed, we rely on the following assumption.
Assumption 2.
Given a random policy $\pi$, assume that
It is worth noting that Assumption 2 holds in MDP problems like grid world navigation (Laskin et al., 2022), which we consider for empirical evaluation in the Experiments section of this paper. Specifically, the action derived from maximizing the return under the random policy becomes (nearly) equivalent to that guided by maximizing the return under the optimal policy, as the maximal return induced by both policies corresponds to navigating to the goal as quickly as possible. We formalize this in Proposition 1 and provide a theoretical proof and empirical validation (refer to Appendix B.2). Besides, we also acknowledge that Assumption 2 may not hold universally for all MDP problems, and we leave its exploration for future research.
Having introduced Assumption 2, we are in conditions of establishing the trustworthiness of the random policy for MDP, which we formally state in the next theorem. To proceed, let us define
$Q_{\text{MDP}}^{\pi,N}(s_{q},a)={\mathbb{E}}_{\pi}\!\!\left[\sum_{t=0}^{N}\gamma^{t}r(s_{t},a_{t})|s_{0}=s_{q},a_{0}=a\right],$ (9)
$\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a)\!=\!\frac{1}{N_{\text{ep}}}\sum_{i=1}^{N_{\text{ep}}}\sum_{t=0}^{N}\!\left(\gamma^{t}r(s_{t},a_{t})|s_{0}=s_{q},a_{0}=a,\pi\right).$ (10)

### v3

@@ 1 Introduction
Pretrained foundation models (FMs) have demonstrated promising performance across a wide variety of domains in artificial intelligence including natural language processing (NLP) (Devlin, 2018; Radford, 2018; Radford et al., 2019; Brown, 2020), computer vision (CV) (Yuan et al., 2021; Sammani et al., 2022; Ma et al., 2023; Chen et al., 2024b), and sequential decision-making (Chen et al., 2021; Janner et al., 2021; Xu et al., 2022b; Yang et al., 2023; Light et al., 2024a; b). This success is attributed to FMs’ impressive capability of in-context learning (Dong et al., 2022; Li et al., 2023; Wei et al., 2023; Wies et al., 2024) which refers to the ability to infer and understand the new (unseen) tasks provided with the context information (or prompt) and without model parameters updates. Recently, in-context reinforcement learning (ICRL) (Laskin et al., 2022; Grigsby et al., 2023; Lin et al., 2023; Sinii et al., 2023; Zisman et al., 2023; Lee et al., 2024; Lu et al., 2024; Wang et al., 2024; Dong et al., 2024) has emerged when FMs are pretrained on sequential decision-making problems. Whereas FMs use texts as the context/prompt in NLP, ICRL treats the state-action-reward tuples as the contextual information for decision-making.
@@
However, the current state-of-the-art (SOTA) ICRL algorithms impose strict requirements on the pretraining datasets. More specifically, Algorithm Distillation (AD) (Laskin et al., 2022) requires the context to contain the complete learning history (from the initial policy to the final-trained policy) of the source (or behavior) RL algorithm for all pretraining environments. In addition, AD requires environments to have short episodes, allowing the context to capture cross-episodic information. This enables AD to learn the improvement operator of the source RL algorithm. Conversely, Decision Pretrained Transformer (DPT) (Lee et al., 2024) partially relaxes the requirement on the context, permitting it to be gathered by random policies and without needing to adhere to the transition dynamics. Nevertheless, DPT necessitates the optimal policy to label an optimal action for any randomly sampled query state across all pretraining environments. To explore the feasibility of ICRL in the absence of optimal policies, Decision Importance Transformer (DIT) (Dong et al., 2024) proposes to leverage the observed state-action pairs in the context data as query states and corresponding action labels. Each state-action pair within the context is assigned a weight in the training process. This weight is proportional to the return-to-go of the pair. Thus, DIT prioritizes the training on high-return pairs. Despite not demanding optimal policies, DIT still requires a substantial context dataset to comprehensively cover all state-action pairs from the state and action spaces. Furthermore, DIT still mandates that more than $30\%$ of the context data comes from well-trained policies to ensure the coverage of good action labels, and the context should originate from a complete episode.
@@
Notably, acquiring either optimal policies or well-trained policies across a multitude of pretraining environments in real-world scenarios can be prohibitively expensive or even intractable. On the other hand, the transition data available in real-world problems—collected as the context—may not originate from a complete episode. These stringent requirements on the pretraining dataset of the SOTA ICRL algorithms severely limit their practical applications to the real world, especially for those where the transition data exhibits high variance to train an effective policy. This work thus aims to relax these requirements by considering an untrained random policy. Although the random policy cannot solve all decision-making problems, it is worth noting that the random policy and the optimal policy select the same optimal actions in certain problems, e.g., the grid world navigation problem (Laskin et al., 2022; Lee et al., 2024; Dong et al., 2024) where the reward is received only upon reaching the unique goal. In these problems, the optimal action induced by both policies corresponds to navigating to the goal as quickly as possible. Consequently, this paper centers on the ICRL that operates without the need for optimal (or any degree of well-trained) policies or episodic context, placing its emphasis on the scenarios under (e.g., uniform) random policies and random contexts only. Our applicable domains in this work focus on the multi-armed bandits and the grid world navigation problems with a single sparse reward received only upon achieving the unique goal.
Figure 1: Schematic of the State-Action Distillation approach: i) Collecting the context by using the random policy to interact with pretraining environments. ii) Sampling a query state randomly from the state space. iii) Starting from the query state and any action in action space, running trust horizons under the random policy, and distilling the action label by the action that yields the maximal return. iv) Pretraining foundation models in a supervised mechanism, which predicts the action label given the context and query state.
@@ 1.1 Main Contributions
2 Related Work
2.1 Offline Reinforcement Learning
2.2 Autoregressive-Supervised Decision Making
2.3 In-Context Reinforcement Learning
@@ 3.1 Preliminaries
RL problems are generally formulated as Markov Decision Processes (MDPs) (Sutton, 2018). An MDP can be represented by a tuple $\tau=(S,A,R,P,\rho)$, where $S$ and $A$ denote finite state and action spaces, $R:S\times A\to\mathbb{R}$ denotes the reward function that evaluates the quality of the action, $P:S\times A\times S\to[0,1]$ denotes the transition probability that describes the dynamics of the system, and $\rho:S\to[0,1]$ denotes the initial state distribution.
@@ 3.2 Supervised Pretraining Mechanism
In this subsection, we introduce the methodology behind ICRL–a supervised pretraining mechanism. Consider two distributions over environments ${\mathcal{T}}_{\text{train}}$ and ${\mathcal{T}}_{\text{test}}$ for pretraining and test (evaluation), respectively. Each environment, along with its corresponding MDP $\tau$, can be regarded as an instance drawn from the environment distributions, where each environment may exhibit distinct reward functions and transition dynamics. Given an environment $\tau$, a context/prompt ${\mathcal{C}}=\{s_{i},a_{i},r_{i},s_{i}^{\prime}\}_{i\in[n]}$ refers to a sample from a pretraining context dataset $D_{\text{train}}(\cdot\mid\tau)$, i.e., ${\mathcal{C}}\sim D_{\text{train}}(\cdot\mid\tau)$, which are collected through interactions between a behavior policy and the environment $\tau$ (see e.g., Algorithm 1). Notably, $D_{\text{train}}(\cdot\mid\tau)$ contains the contextual information regarding the environment $\tau$. We next consider a query state distribution $D_{q}^{\tau}$ and a label policy that maps the query state to the distribution of the action label, i.e., $\pi_{l}:S\to\Delta_{a_{l}}(A)$, where $\Delta_{a_{l}}(\cdot)$ denotes the probability distribution and $a_{l}$ denotes the prediction/output goal of the FM. Then, the joint distribution over the environment $\tau$, context ${\mathcal{C}}$, query state $s_{q}$, and action label $a_{l}$ is given by
@@
where $l(\cdot,\cdot)$ represents the loss function, for instance, negative log-likelihood (NLL) for discrete-action problems and mean square error (MSE) for continuous-action scenarios. We note that while the current SOTA ICRL algorithms (AD, DPT, and DIT) adhere to a common objective function equation 3, they differ significantly in constructing the context, query state, and action label. Furthermore, it is important to highlight that selecting appropriate action labels in existing ICRL algorithms can be prohibitively expensive. We discuss this in more detail for each of the ICRL baselines that follow.
@@
Instead of learning an optimal policy for a specific environment, AD proposes to learn a RL algorithm itself across a wide range of environments. This is, the FM in AD is pretrained to imitate the source (or behavior) algorithm over the pretraining environment distribution ${\mathcal{T}}_{\text{train}}$. In general, AD demands a well-trained source algorithm with its complete learning history (from the initial policy to the final-trained policy). Additionally, AD is restricted to the environments with short episode length, as the context must capture cross-episode information of the source algorithm, while standard transformers often have limited context length. In terms of the objective function equation 3, AD takes the state at the time step $t$ as the query state, $a_{t}$ from the source algorithm as the action label, and the episodic history data $(s_{0},a_{0},r_{0},\cdots,s_{t-1},a_{t-1},r_{t-1})$ to be the context.
@@
Instead of being stringent on the context and the environment itself, DPT handles the context and query state in a more general manner. Specifically, in terms of the objective function equation 3, DPT can consider a random collection of transitions as the context ${\mathcal{C}}$, a random query state $s_{q}$ drawn from $D_{q}$, and an optimal action label corresponding to $s_{q}$. Despite less requirements on the context, DPT necessitates access to optimal policies for the optimal action labels in all pretraining environments, which may not be available in real-world applications.
@@
DIT proposes to learn ICRL without optimal action labels, following the same supervised pretraining mechanism as DPT. To that end, DIT chooses to consider every possible state and action in the context as the query state and corresponding action label. It is important to point out that DIT requires a (partially) complete episode to form the context, enabling the computation of a return-to-go for each state-action pair. By mapping the return-to-go to a weight that reflects the quality of each state-action pair, DIT can pretrain the FM using the DPT structure, augmented by the weight assigned to each query state and action label. In other words, DIT prioritizes the training on high-return pairs. Notably, DIT still mandates that more than $30\%$ of the context data comes from well-trained policies to ensure the coverage of good action labels.
To summarize, it is worth highlighting that all these SOTA ICRL algorithms necessitate varying degrees of well-trained, or even optimal policies during the pretraining phase. However, obtaining such policies for real-world applications is often prohibitive, as it demands extensive training across a vast number of real-world environments. This challenge becomes even more pronounced in the domains where the transition data exhibits high variance to train an effective policy. Thus, executing ICRL under (e.g., uniform) random policies and random contexts is crucial for enabling the practical application of ICRL in the real-world.
@@ 4 State-Action Distillation
In this section, we propose the State-Action Distillation (SAD), an approach for generating the pretraining dataset for ICRL under random policies and random contexts (see Figure 1).
As indicated in equation 2, the pretraining data consists of the context, the query state and the action label. We start by introducing the generation of the context under random policies in SAD (refer to Algorithm 1). It is important to highlight that the context is collected through interactions with the environment under any given random policy (e.g., uniform random policy), and notably, the context not necessarily originates from a complete episode. These benefits make SAD potentially well-suited for ICRL’s real-world applications with random transition data only.
@@
We proceed by recalling that DIT prioritizes the training on high-return pairs from the context data collected. To that end, DIT assigns a weight $w$ to the loss function during the pretraining phase that is proportional to the return-to-go, i.e., $w(s_{t},a_{t})\propto\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}.$ Nonetheless, we acknowledge that DIT may not explore to train on good state-action pairs under the random policy for two reasons: ($i$) DIT solely considers to train on the state-action pairs that are observed in the context, which now contains limited and, more importantly, random transition data. ($ii$) Even for the state-action pairs in the collected context, the return-to-go does not necessarily prioritize the optimal pair but rather promotes the pair with high immediate reward, as the discount factor applies starting from the current time step with a horizon of $(T-t+1)$ only, instead of $T$. This issue becomes even more critical in the problems with sparse rewards.
Under this circumstance, our SAD approach advocates for distilling the outstanding query states and action labels by searching across the entire state and action spaces under the random policy. Before proceeding, we recall the definition of the optimal action in the problems of multi-armed bandit (MAB) and MDP. For any query state $s_{q}$, the optimal action for $s_{q}$ corresponds to the action that maximizes the optimal Q-function
$a_{\text{MAB}}^{*}(s_{q})\overset{\Delta}{=}\operatorname*{arg\,max}_{a\in A}\underbrace{{\mathbb{E}}\left[r(s_{q},a)\right]}_{Q_{\text{MAB}}^{*}(s_{q},a)},\quad a_{\text{MDP}}^{*}(s_{q})\overset{\Delta}{=}\operatorname*{arg\,max}_{a\in A}\underbrace{{\mathbb{E}}_{\pi^{*}}\!\!\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}=s_{q},a_{0}=a\right]}_{Q_{\text{MDP}}^{*}(s_{q},a)},$ (4)
where $s_{q}$ in the MAB problem refers to the singleton state of bandits, and $\pi^{*}$ denotes the optimal policy in the MDP. However, both $a_{\text{MAB}}^{*}$ and $a_{\text{MDP}}^{*}$ are intractable to obtain in our problem of interest for two reasons: ($i$) computing the expectation in the Q-function demands to sample infinite episodes; ($ii$) one can only have access to the random policy, instead of $\pi^{*}$. Therefore, we instead consider ($i$) stochastic approximation that uses the average as the unbiased estimate of the expectation due to the law of large numbers; ($ii$) maximizing the episodic return under the random policy.
@@ 4.1 Trustworthiness of the Random Policy
Subsequently, the crucial question arises: when can we trust the random policy? We claim: The random policy is probabilistically trustworthy for the MAB and MDP problems within a trust horizon. We formalize this claim for MAB and MDP in this subsection, which relies on the following assumptions.
@@
Note that Assumption 1 is common in the literature (Azar et al., 2017; Wei et al., 2020; Zhang et al., 2021). In particular, in the case of finite state-action spaces, it is always possible to design the reward to avoid the possibility of being unbounded.
Assumption 2.
Given a random policy $\pi$, assume that
It is worth highlighting first that Assumption 2 may not hold universally for all MDP problems. In fact, it is unlikely that a random policy and the optimal policy select the same optimal actions for all problems. Nevertheless, we acknowledge that Assumption 2 does hold in the MDP problems like the grid world navigation where the reward is received only upon achieving the unique goal (Laskin et al., 2022; Lee et al., 2024; Dong et al., 2024), which fall into the applicable domains of this work. We visualize in Figure 2 a single-dimensional grid world navigation problem as an example. In particular, the action derived from maximizing the return under the random policy becomes equivalent to that guided by maximizing the return under the optimal policy, as the maximal return induced by both policies corresponds to navigating to the goal as quickly as possible. We formalize this in Proposition 1 and provide a theoretical proof and empirical validation (refer to Appendix A.2). Although Assumption 2 holds for the grid world navigation problems considered in this work, it is crucial to point out that Assumption 2 may not hold for all grid world navigation problems. To this end, we further discuss the validity of Assumption 2 for the grid world navigation with two sparse rewards (see Appendix A.3), and demonstrate that the validity of Assumption 2 depends on problem-specific factors such as the rewards and the discount factor $\gamma$.
Figure 2: A single-dimensional grid world MDP comprising five states $\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$, where $s_{0}$ represents the goal state (golden star). The environment offers two possible actions: $a_{0}$ (go left), and $a_{1}$ (go right). Any transitions that would result in (left or right) boundary crossing will be confined to the current position. The reward structure is sparse, with a value of 1 received solely upon reaching the unique goal state $s_{0}$ and a value of 0 otherwise. We consider an infinite time horizon with a discounter factor $\gamma$.
Having introduced the necessary assumptions, we are in the stage of investigating the trustworthiness of the random policy in both MAB and MDP problems. To proceed, we rely on the following two definitions.
Definition 1 (MAB).
Denote by $s_{q}$ and $a^{*}$ the singleton state and optimal arm in the MAB problem. Consider a random policy $\pi$. Suppose that each arm $a$ has been selected $N_{a}$ times $(N_{a}\in{\mathbb{N}}_{+})$ under $\pi$. The trustworthiness of the random policy in the the MAB problems represents the probability of selecting the optimal arm under the random policy, i.e.,
$\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r_{i}(s_{q},a^{*})\geq\max_{a\in A\setminus\{a^{*}\}}\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r_{i}(s_{q},a).$ (6)
The corresponding trust horizon denotes the minimal number of selections over all arms, i.e., $N\triangleq\min_{a\in A}N_{a}$ that guarantees such trustworthiness.
Definition 1 implies by the law of large numbers that $N\to\infty$ serves as a trust horizon, ensuring the selection of the optimal arm with probability one. In practice, achieving large trustworthiness requires a sufficiently large trust horizon. The intuition is that a large enough horizon approximates the infinite step problem. The same intuition holds for the MDP problems. We formalize this concept in the next definition, which builds upon the Q-function of the finite-horizon MDP
$Q_{\text{MDP}}^{\pi,N}(s_{q},a)={\mathbb{E}}_{\pi}\!\!\left[\sum_{t=0}^{N}\gamma^{t}r(s_{t},a_{t})|s_{0}=s_{q},a_{0}=a\right],\,\forall a\in A,$ (7)
$\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a)\!=\!\frac{1}{N_{\text{ep}}}\sum_{i=1}^{N_{\text{ep}}}\sum_{t=0}^{N}\!\left(\gamma^{t}r(s_{t},a_{t})|s_{0}=s_{q},a_{0}=a,\pi\right),\,\forall a\in A,$ (8)
Definition 2 (MDP).
Denote by $a^{*}$ the optimal action given a query state $s_{q}$. Consider a random policy $\pi$ as well as its Q-function $Q_{\text{MDP}}^{\pi}$. The trustworthiness of the random policy in the the MDP problems represents the probability of selecting the optimal action for any query state in the state space, i.e.,
$\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a^{*})\geq\max_{a\in A\setminus\{a^{*}\}}\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a),\,\forall s_{q}\in S.$ (9)
The corresponding trust horizon denotes the minimal horizon length of the MDP such that
$Q_{\text{MDP}}^{\pi,N}(s_{q},a^{*})\geq\max_{a\in A\setminus\{a^{*}\}}Q_{\text{MDP}}^{\pi,N}(s_{q},a),\,\forall s_{q}\in S.$ (10)
In both Definitions 1 and 2, the trustworthiness refers to the probability of selecting the optimal action. While the trust horizon in both cases approximates the infinite-horizon problem, they show subtle differences in the relationship to the trustworthiness. In the MAB problem, which consists of a single episode, the trustworthiness depends on the trust horizon only. However, in the MDP, computing $Q_{\text{MDP}}^{\pi,N}(s_{q},a)$ is generally intractable, thus requiring the use of an unbiased estimate (as defined in equation 8), which depends on the number of episodes $N_{\text{ep}}$. Consequently, the trustworthiness in the MDP setting is influenced by both the trust horizon and the number of episodes. Next, we formally quantify the relationship between the trustworthiness and the trust horizon for the MAB and MDP problems.
@@
Let Assumption 1 hold. The random policy is at least $(1-\delta)$-trustworthy as in Definition 1, when the trust horizon $N$ satisfies
@@
Proof. See Appendix A.1. $\square$
Theorem 1 implies that the trust horizon $N$ quantifies the trustworthiness of the decision making under the random policy $\pi$ for MAB problems. Indeed, a larger $N$ implies a higher probability (smaller $\delta$) that the average reward of the optimal arm under $\pi$ exceeds that of the next-best arm, therefore, making a more reliable decision. We substantiate this claim by empirical evidence (depicted in Figure 6(a)). In the practical implementation, we simply execute the MAB under the random policy $\pi$ until every action in the action space $A$ selected at least $N$ times. Subsequently, we select the action with the maximal average reward as the action label. The detailed procedure for collecting such action labels in MAB is outlined in Algorithm 2.
@@
2:  Execute the MAB in $\tau$ under the random policy $\pi$ until every action in $A$ selected at least $N$ times